OpenStudy (anonymous):

Need help ! :/ Liam deposited $1,724.53 in a savings account 4 months ago. The account earns 3.1% simple interest. What is his account balance today? $1,724.53 $1,742.35 $1,777.99 $2,205.67

OpenStudy (anonymous):

http://www.webmath.com/_answer.php

OpenStudy (jdoe0001):

rate => 3.1% = 3.1/100 = 0.031 principal => original amount periods => 4 months, thus = 4

OpenStudy (jdoe0001):

use the formula for "simple interest" and get the interest it earned thus far ADD that interest earned, to the original amount and that's your new amount, or so-called "balance"

OpenStudy (anonymous):

\[1724.53\left(1+4\left(\frac{.031}{12}\right)\right)=1742.35 \] Sorry, the prior posting had an incorrect answer.